# SafeInfer: Context Adaptive Decoding Time Safety Alignment for Large Language Models

### v1

@@ Abstract
Safety-aligned language models often exhibit fragile and imbalanced safety mechanisms, increasing the likelihood of generating unsafe content. In addition, incorporating new knowledge through editing techniques to language models can further compromise safety. To address these issues, we propose SafeInfer, a context-adaptive, decoding-time safety alignment strategy for generating safe responses to user queries. SafeInfer comprises two phases: the ‘safety amplification’ phase, which employs safe demonstration examples to adjust the model’s hidden states and increase the likelihood of safer outputs, and the ‘safety-guided decoding’ phase, which influences token selection based on safety-optimized distributions, ensuring the generated content complies with ethical guidelines. Further, we present HarmEval, a novel benchmark for extensive safety evaluations, designed to address potential misuse scenarios in accordance with the policies of leading AI tech giants. We release the source code and dataset at: https://github.com/NeuralSentinel/SafeInfer.
[…]
[ topline=false, bottomline=false, skipabove=skipbelow=leftline=true, rightline=false, linecolor=red, linewidth=4pt, innertopmargin=2pt, innerbottommargin=2pt, innerrightmargin=2pt, innerleftmargin=2pt, backgroundcolor=gray!10, roundcorner=10pt ]stylishframe
SafeInfer: Context Adaptive Decoding Time
Safety Alignment for Large Language Models
[…]
1 Introduction
The extensive use of LLMs in various applications presents substantial challenges in safety and ethical alignment Weidinger et al. (2021); Wang et al. (2023), particularly in environments that demand strict adherence to ethical standards.
@@ 1 Introduction
Among the prominent issues is ‘jailbreaking’, where models circumvent built-in restrictions to generate undesirable content Banerjee et al. (2024); Deng et al. (2024); Zou et al. (2023b), thereby exposing the limitations of traditional prompting methods that may inadvertently trigger sensitive topics. Traditional fine-tuning offers a measure of control by retraining models on specific datasets, but it falls short in effectively managing complex inputs that can provoke such issues Qi et al. (2024). Instead, decoding time alignment, through techniques like controlled text generation (CTG) Liu et al. (2021a), offers a more nuanced solution by allowing dynamic, real-time moderation of outputs without necessitating changes to the model’s architecture or extensive retraining. This approach tailors outputs directly in response to the input context, ensuring certain attribute (such as detoxification, politeness) aligned interactions across various applications Huang et al. (2024). In parallel, previous studies Subramani et al. (2022); Hernandez et al. (2023); Zou et al. (2023a); Todd et al. (2024) have demonstrated that the in-context learning mechanism can guide specific tasks through the model’s activations. Activation engineering techniques have shown promise in steering model behavior by manipulating these activations.
Drawing on these findings, we introduce SafeInfer, an novel strategy for in-context adaptive decoding time alignment which comprises two phases, as illustrated in Figure 1. The initial phase, termed as Safety amplification (SA) phase, utilizes demonstration examples to derive the safety amplification vector, which is then integrated into the hidden state of the language model. The second phase employs a Safety guided decoding strategy (sGDS) that combines/removes the biased attributes through the integration of different distributions from language models. This phase enhances safety by preferentially selecting tokens from certain distributions over others, thereby optimizing the overall output distribution for safety. In this paper, our primary objective is to realign the model toward heightened safety by employing contextual adaptation alongside a decoding strategy. This approach not only prioritizes safety alignment but also ensures the preservation of the overall utility benchmark of the language model. In addition, we have designed this methodology to be seamlessly adaptable to different language model architectures, thereby broadening its utility and applicability in a variety of settings.
@@
  • To best of our knowledge, we are the first to apply our strategy across both the base and edited versions of widely used large language models, evaluating them on six distinct datasets.
@@ 2 Related work
Inference time safety alignment: Ensuring the safety and robustness of AI models without retraining involves several approaches. Training-free methods like rule-based filtering Feng et al. (2020) and ensemble techniques enhance safety by filtering harmful or biased content and using multiple models to cross-verify outputs Liang et al. (2023); Lu et al. (2022); Qin et al. (2022). Decoding-time safety alignment modifies the generation process with constrained decoding to prioritize safe outputs Gehman et al. (2020); Dathathri et al. (2020); Wan et al. (2023); Huang et al. (2024). Inference-time safety alignment focuses on real-time monitoring and intervention, using reinforcement learning from human feedback (RLHF) to adjust model behavior based on feedback Ouyang et al. (2022) and adversarial training to improve robustness. Recent work explores modular approaches like Bai et al. (2022); Xu et al. (2024).
Controlled text generation: Techniques for CTG steer the outputs of a language model to align with specific attributes like style. This is achieved by modifying the model’s output probabilities, typically using a parameter that determines the degree of this modulation. Strategies include using dedicated classifiers Yang and Klein (2021); Sansone and Manhaeve (2023); Kim et al. (2023), specially fine-tuned smaller models Liu et al. (2021b), or varying the prompts fed into the same language model Pei et al. (2023); Sanchez et al. (2024). Many CTG methods apply concepts akin to those in Bayes’ theorem to effectively skew the model’s responses toward the intended attributes Hallinan et al. (2023).
3 SafeInfer: Context Adaptive Decoding Time Safety Alignment
@@
Safety amplification (SA): This phase is designed to control the latent space of the target model $M_{t}$ by leading it through the safety guided demonstrations $D_{sf}$. Following the approach described in Todd et al. (2024) for encoding task-specific guided demonstrations into a vectorized form, we obtain the $\mathsf{SV}$ using the dataset $D_{sf}$. Further, the $\mathsf{SV}$ is integrated at certain layer during the forward pass through $M_{t}$. The detailed process is explained in the subsequent paragraph.
Computing safety amplification vector ($\mathsf{SV}$): This computation involves identifying top attention heads through activation patching Zhang and Nanda (2024); Todd et al. (2024); Makelov et al. (2024), preparing prompt from $D_{sf}$ and obtaining safety amplification vector $\mathsf{SV}$. For identifying influential heads in language model, we solely follow the approach provided by Todd et al. (2024) (see Appendix A.1). We denote the set of influential attention heads as $A$, where each attention head at layer $l$ and position $j$ is represented by $attn_{lj}$. From $D_{sf}$, we construct a set of prompts $\mathsf{P}$, where each prompt $\mathsf{p}\in\mathsf{P}$ is structured as $\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n}),q_{n+1}\}$. For each attention head $attn_{lj}$, we compute the mean of the representations of the prompts $\mathsf{P}$ and denote it as safety conditioned activations $attn_{lj}^{{}^{\prime}}$, as shown in Equation 1.
@@
We incorporate the $\mathsf{SV}$ into the hidden state ($h_{l}$) of the target model $M_{t}$ at layer $l$ to perform safety amplification (Equation 3), thereby obtaining the updated hidden state $h_{l}^{{}^{\prime}}$. We follow Todd et al. (2024) for selecting the layer $l$. We denote the target model with the updated hidden state as $M_{t}^{{}^{\prime}}$. The coefficient $\gamma$ is a hyperparameter.
@@
Safety guided decoding strategy (sGDS): In this phase, we aim to further enhance the safety of the model $M_{t}^{{}^{\prime}}$ by controlling the next token generation during the decoding process. The intention is to mitigate certain negative attributes, such as harm and unethical behavior, by debiasing the output distribution of $M_{t}^{{}^{\prime}}$. We begin by fine-tuning a language model of same family as $M_{b}$ using a dataset $\mathbb{D}_{usf}$, resulting in the model $M_{usf}$. This model inherently exhibits a bias toward generating harmful responses. For example, it is more likely to predict the word “Sure” rather than “Sorry” as the initial token in response to a harmful query. To achieve safe and helpful generation, it is crucial to preserve the original distribution of $M_{t}^{{}^{\prime}}$ while mitigating the harmful tendencies observed in $M_{usf}$. This requires addressing such harmful tendencies without significantly altering the overall behavior or output distribution of $M_{t}^{{}^{\prime}}$. To accomplish this, we employ CTG strategy proposed in  Dekoninck et al. (2023). We first obtain a combined distribution $\mathscr{C}$ that integrates the output distributions of both $M_{t}^{{}^{\prime}}$ and $M_{usf}$, allowing for distinct attributes (e.g., harms, biases) while preserving abilities from both distributions. We use Union operation Dekoninck et al. (2023) to obtain the distribution $\mathscr{C}$. This operator enables a non-linear combination of the two distributions $M_{t}^{{}^{\prime}}$ and $M_{usf}$, such that if either $M_{t}^{{}^{\prime}}$ or $M_{usf}$ assigns a high probability to a particular token $x$, the resulting distribution will reflect a similarly high probability for that token. The optimization function, based on Kullback-Leibler divergence, is provided in Equation 4, where $I(x)$ is the indicator function.
@@
Following Dekoninck et al. (2023), we obtain the distribution $\mathscr{C}$ using the solution of the optimization function presented in Equation 5. $\sigma$ denotes the standard softmax.
@@
$\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}M_{t}^{sf}$ $=M_{t}^{{}^{\prime}}-\lambda\cdot\sigma(\max(M_{t}^{{}^{\prime}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{usf}}))$
@@
4 Datasets
We evaluate SafeInfer on five existing datasets – DangerousQA Shaikh et al. (2023), AdvBench Zou et al. (2023b), HEx-PHI Qi et al. (2023a), NicheHazardQA Hazra et al. (2024), and TechHazardQA Banerjee et al. (2024). Further, we propose a new safety dataset based on the list of violated policies identified by Meta Qi et al. (2023a). We describe each of these datasets in detail below.
@@ 4 Datasets
HarmEval: To further evaluate our safety strategy SafeInfer across a broad spectrum of potential hazards, we developed HarmEval, a dataset based on prohibited scenarios listed in OpenAI and Meta’s usage policies, as cited in Qi et al. (2023b). HarmEval categorizes risks into 11 main categories (see Figure 3), resulting in approximately $\sim$550 crafted harmful queries. We employed a two-step verification process for these queries. First, we used GPT-4 to classify the queries as harmful or non-harmful [^2] Prompt: Determine whether the question adheres to ethical standards or if there are certain contents that raise ethical concerns. Please answer only in either ‘Yes’ or ‘No’.. Following this, we utilized the Perspective API [^3] https://perspectiveapi.com/ to calculate the toxicity score, selecting only those queries with a toxicity score of 0.7 or higher. The dataset will be released upon the acceptance of our work.
@@
Llama-2 Mistral Llama-2 Mistral Llama-2 Mistral
Base model 17.83 5.22 46.90 62.00 0.298, 0.451 0.501, 0.656
SafeDecoding 80.30 - 45.70 - 0.376, 0.518 -
SafeInfer 20.09 5.22 46.47 61.60 0.390, 0.582 0.531, 0.691
@@ 5.1 Language models
Base models: In accordance with Jain et al. (2023), we utilize base model backbones such as Llama2-7b-chat-hf Touvron et al. (2023) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023).
Edited models: Previous research Banerjee et al. (2024); Hazra et al. (2024) has observed that edited models can introduce hidden harms after updating the knowledge of the model (model editing). Therefore, our method has been evaluated on edited models with the Llama2-7b-chat-hf backbone. We employ a locate-and-edit model-based algorithm known as ROME Meng et al. (2022). Our primary goal is to examine the impact of model editing on model safety, which is why we opted for a single edit algorithm (ROME) and a single model (Llama-2). For the most part, we utilize the default parameter values provided in paper Hazra et al. (2024).
5.2 Prompting technique
@@
For simple prompts, we employed the vanilla strategy by directly asking the questions present in the datasets and expecting the model to generate responses. Recent studies by Banerjee et al. (2024) have demonstrated that models can be ‘jailbroken’ by prompting them in an instruction-centric manner. This is followed by instruction-centric CoT prompts, which infuse unethical content more effectively into the generated responses. Inspired by this, we conduct experiment using instruction-centric and instruction-centric CoT prompts.
To assess the defense performance when a naive attacker directly inputs harmful queries to the language model, we utilized the six datasets mentioned previously. Detailed setups of these prompting techniques can be found in the Appendix A.3.
5.3 Baselines
We evaluate our proposed method against the following safety alignment baselines following a decoding-based approach: SafeDecoding Xu et al. (2024) and Self-CD Shi et al. (2024) methods. Further, we directly use SA and sGDS as a standalone baselines to establish the effectiveness of the amalgamation of the two techniques.
SafeDecoding: SafeDecoding Xu et al. (2024) is a safety decoding strategy used while responding to user queries. This approach is built upon the crucial observation that tokens representing safety warnings are often ranked high in probability, even when harmful content tokens are also prevalent. By selectively boosting the probability of these safety tokens and diminishing the likelihood of harmful sequences, SafeDecoding effectively counters the risks posed by jailbreak attacks. We show the results of the Llama2-7b model. Due to the lack of knowledge about the fine-tuning dataset used, we could not reproduce the results for Mistral-7b.
Self-CD: We also compare SafeInfer against Self-Contrastive Decoding (Self-CD) Shi et al. (2024), which mitigates the issues of harmfulness as well as helpfulness. Self-CD is designed as a training-free and model-independent intervention, which attempts to amplify the difference in output token distributions when responding to questions with a safety prompt and without a safety prompt. The final next token distribution is determined by removing the over-attention from the model via contrastive decoding.
@@ 5.3 Baselines
5.4 Evaluation metric
We follow the methodology outlined by Liu et al. (2024) and utilize attack success rate (ASR) to evaluate the effectiveness of SafeInfer. ASR is defined as follows.
@@ 5.4 Evaluation metric
Simulating GPT-4 as an evaluator: Manually evaluating each generated response for potential harm is both tedious and mentally taxing. To address this, inspired by several prior works Qi et al. (2023b); Zheng et al. (2023), we employ GPT-4 as an automated evaluator to simulate human judges. We instruct GPT-4 to analyze and assess the ethical implications of responses generated by our test models. The specific instruction given to GPT-4 is: “Given the following text, analyze and assess its ethical implications. Determine whether the content adheres to ethical standards or if there are elements that raise ethical concerns.” GPT-4 is then tasked with providing a ‘YES’ or ‘NO’ answer regarding the presence of ethical concerns. To ensure the accuracy of GPT-4’s evaluations, we randomly sample 40% of the model-generated responses and obtain human judgments for comparison. Four experts with experience in working on AI safety partook in the evaluation. The average pairwise inter-annotator agreement (Cohen’s $\kappa$) is 0.94. Our analysis reveals a high concordance rate, with GPT-4’s judgments matching human judgments in $>93\%$ cases. Note that to manage computational costs, we select a stratified sample of up to 100 questions from of the each categories present in the NicheHazardQA, TechHazardQA, and HarmEval datasets. When fewer than 100 questions were available in a category, we use all available questions. We average the results from over all the categories. For other datasets – DangerousQA, AdvBench, and HEx-PHI – we select $\sim$200 stratified questions. For every dataset the selected questions are fed to the model, and the responses are evaluated for safety using GPT-4 and humans.
5.5 Obtaining the harmful model
We construct a small set of safe demonstrations, $D_{sf}$, from our proposed HarmEval dataset, consisting of approximately $|\mathsf{P}|$ = 100 prompts. Each prompt, $\mathsf{p}$, includes 10 contextual samples (harmful question-safe answer (see samples in Appendix A.4)) and a query. Further, we use the HarmEval dataset to create $\mathbb{D}_{usf}$, a collection of harmful question-answer pairs. Following Qi et al. (2023a), we select around $\sim$100 queries and their harmful responses to finetune a model with the same base model as $M_{b}$ and obtain the harmful model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{usf}$.
5.6 Utility and over-safety test
To evaluate the utility of the model after applying the proposed method, we conduct thorough evaluation on MMLU (5 shots) Hendrycks et al. (2021) and TruthfulQA Lin et al. (2022). For testing over-safety, we use the framework used by Röttger et al. (2024) where the LLM backbone generates three main types of responses on the XSTest dataset: (1) full compliance (2) full refusal (3) partial refusal. We only count responses classified as full compliance as the refusal rate to measure over-safety.
6 Results

### v2

@@ Abstract
Language models aligned for safety often exhibit fragile and imbalanced mechanisms, increasing the chances of producing unsafe content. In addition, editing techniques to incorporate new knowledge can further compromise safety. To tackle these issues, we propose SafeInfer, a context-adaptive, decoding-time safety alignment strategy for generating safe responses to user queries. SafeInfer involves two phases: the ‘safety amplification’ phase, which uses safe demonstration examples to adjust the model’s hidden states and increase the likelihood of safer outputs, and the ‘safety-guided decoding’ phase, which influences token selection based on safety-optimized distributions to ensure the generated content adheres to ethical guidelines. Further, we introduce HarmEval, a novel benchmark for comprehensive safety evaluations, designed to address potential misuse scenarios in line with the policies of leading AI technology companies. We release the source code and dataset at: https://github.com/NeuralSentinel/SafeInfer.
Introduction
The extensive use of LLMs in various applications presents substantial challenges in safety and ethical alignment (Weidinger et al. 2021; Wang et al. 2023), particularly in environments that demand strict adherence to ethical standards. Among the prominent issues is ‘jailbreaking’, where models circumvent built-in restrictions to generate undesirable content (Banerjee et al. 2024; Deng et al. 2024; Zou et al. 2023b), thereby exposing the limitations of traditional prompting methods that may inadvertently trigger sensitive topics. Traditional fine-tuning offers a measure of control by retraining models on specific datasets, but it falls short in effectively managing complex inputs that can provoke such issues (Qi et al. 2024). Instead, decoding time alignment, through techniques like controlled text generation (CTG) (Liu et al. 2021a), offers a more nuanced solution by allowing dynamic, real-time moderation of outputs without necessitating changes to the model’s architecture or extensive retraining. This approach tailors outputs directly in response to the input context, ensuring certain attribute (such as detoxification, politeness) aligned interactions across various applications (Huang et al. 2024).
@@ Introduction
In parallel, previous studies (Subramani, Suresh, and Peters 2022; Hernandez, Li, and Andreas 2023; Zou et al. 2023a; Todd et al. 2024) have demonstrated that the in-context learning mechanism can guide specific tasks through the model’s activations. Activation engineering techniques have shown promise in steering model behavior by manipulating these activations.
Drawing on these findings, we introduce SafeInfer, an novel strategy for in-context adaptive decoding time alignment which comprises two phases, as illustrated in Figure 1. The initial phase, termed as Safety amplification (SA) phase, utilizes demonstration examples to derive the safety amplification vector, which is then integrated into the hidden state of the language model. The second phase employs a Safety guided decoding strategy (sGDS) that combines/removes the biased attributes through the integration of different distributions from language models. This phase enhances safety by preferentially selecting tokens from certain distributions over others, thereby optimizing the overall output distribution for safety. The key novelty of our work lies in judiciously coupling these two phases to reap benefits from each of them to ensure a more effective safety alignment compared to what is existing in the literature. The first phase is motivated by the recent works which proved that moving the latent space of the model toward a specific task can help the model to actually solve the task better (Todd et al. 2024; Liu et al. 2024a). For the decoding time intervention, we next use the concept of controlled text generation in the lines of (Dekoninck et al. 2024). We do not know of any work that couples these two ideas simultaneously to achieve safety alignment. Overall, in this paper, our primary objective is to realign the model toward heightened safety by employing contextual adaptation alongside a decoding strategy. This approach not only prioritizes safety alignment but also ensures the preservation of the overall utility benchmark of the language model. In addition, we have designed this methodology to be seamlessly adaptable to different language model architectures, thereby broadening its utility and applicability in a variety of settings.
@@
  • To best of our knowledge, we are the first to apply our strategy across both the base and edited versions of widely used large language models, evaluating them on six distinct datasets. We demonstrate that our approach not only drastically reduces the number of harmful responses by SOTA LLMs but is also able to preserve the basic utilities of these LLMs as evidenced by five open-ended benchmark tasks.
@@
Inference time safety alignment: Ensuring the safety and robustness of AI models without retraining involves several approaches. Training-free methods like rule-based filtering (Feng et al. 2020) and ensemble techniques enhance safety by filtering harmful or biased content and using multiple models to cross-verify outputs (Liang et al. 2023; Lu et al. 2022; Qin et al. 2022). Decoding-time safety alignment modifies the generation process with constrained decoding to prioritize safe outputs (Gehman et al. 2020; Dathathri et al. 2020; Wan et al. 2023; Huang et al. 2024). Inference-time safety alignment focuses on real-time monitoring and intervention, using reinforcement learning from human feedback (RLHF) to adjust model behavior based on feedback (Ouyang et al. 2022) and adversarial training to improve robustness. Recent work explores modular approaches like (Bai et al. 2022; Xu et al. 2024).
Controlled text generation: Techniques for CTG steer the outputs of a language model to align with specific attributes like style. This is achieved by modifying the model’s output probabilities, typically using a parameter that determines the degree of this modulation. Strategies include using dedicated classifiers (Yang and Klein 2021; Sansone and Manhaeve 2023; Kim et al. 2023), specially fine-tuned smaller models (Liu et al. 2021b), or varying the prompts fed into the same language model (Pei, Yang, and Klein 2023; Sanchez et al. 2024). Many CTG methods apply concepts akin to those in Bayes’ theorem to effectively skew the model’s responses toward the intended attributes (Hallinan et al. 2023).
SafeInfer: Context Adaptive Decoding Time Safety Alignment
@@
Safety amplification (SA): This phase is designed to control the latent space of the target model $M_{t}$ by leading it through the safety guided demonstrations $D_{sf}$. Following the approach described in (Todd et al. 2024) for encoding task-specific guided demonstrations into a vectorized form, we obtain the $\mathsf{SV}$ using the dataset $D_{sf}$. Further, the $\mathsf{SV}$ is integrated at certain layer during the forward pass through $M_{t}$. The detailed process is explained in the subsequent paragraph.
Computing safety amplification vector ($\mathsf{SV}$): This computation involves identifying top attention heads through activation patching (Zhang and Nanda 2024; Todd et al. 2024; Makelov et al. 2024), preparing prompt from $D_{sf}$ and obtaining safety amplification vector $\mathsf{SV}$. For identifying influential heads in language model, we solely follow the approach provided by (Todd et al. 2024). We denote the set of influential attention heads as $A$, where each attention head at layer $l$ and position $j$ is represented by $attn_{lj}$. From $D_{sf}$, we construct a set of prompts $\mathsf{P}$, where each prompt $\mathsf{p}\in\mathsf{P}$ is structured as $\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n}),q_{n+1}\}$. For each attention head $attn_{lj}$, we compute the mean of the representations of the prompts $\mathsf{P}$ and denote it as safety conditioned activations $attn_{lj}^{{}^{\prime}}$, as shown in Equation 1.
@@
We incorporate the $\mathsf{SV}$ into the hidden state ($h_{l}$) of the target model $M_{t}$ at layer $l$ to perform safety amplification (Equation 3), thereby obtaining the updated hidden state $h_{l}^{{}^{\prime}}$. We follow (Todd et al. 2024) for selecting the layer $l$. We denote the target model with the updated hidden state as $M_{t}^{{}^{\prime}}$. The coefficient $\gamma$ is a hyperparameter.
@@
Safety guided decoding strategy (sGDS): In this phase, we aim to further enhance the safety of the model $M_{t}^{{}^{\prime}}$ by controlling the next token generation during the decoding process. The intention is to mitigate certain negative attributes, such as harm and unethical behavior, by debiasing the output distribution of $M_{t}^{{}^{\prime}}$. We begin by fine-tuning a language model of same family as $M_{b}$ using a dataset $\mathbb{D}_{usf}$, resulting in the model $M_{usf}$. This model inherently exhibits a bias toward generating harmful responses. For example, it is more likely to predict the word “Sure” rather than “Sorry” as the initial token in response to a harmful query. To achieve safe and helpful generation, it is crucial to preserve the original distribution of $M_{t}^{{}^{\prime}}$ while mitigating the harmful tendencies observed in $M_{usf}$. This requires addressing such harmful tendencies without significantly altering the overall behavior or output distribution of $M_{t}^{{}^{\prime}}$. To accomplish this, we employ CTG strategy proposed in  (Dekoninck et al. 2023). We first obtain a combined distribution $\mathscr{C}$ that integrates the output distributions of both $M_{t}^{{}^{\prime}}$ and $M_{usf}$, allowing for distinct attributes (e.g., harms, biases) while preserving abilities from both distributions. We use Union operation (Dekoninck et al. 2023) to obtain the distribution $\mathscr{C}$. This operator enables a non-linear combination of the two distributions $M_{t}^{{}^{\prime}}$ and $M_{usf}$, such that if either $M_{t}^{{}^{\prime}}$ or $M_{usf}$ assigns a high probability to a particular token $x$, the resulting distribution will reflect a similarly high probability for that token. The optimization function, based on Kullback-Leibler divergence, is provided in Equation 4, where $I(x)$ is the indicator function.
@@
Following (Dekoninck et al. 2023), we obtain the distribution $\mathscr{C}$ using the solution of the optimization function presented in Equation 5. $\sigma$ denotes the standard softmax.
@@
$\color[rgb]{0.1333,0.5451,0.1333}\definecolor[named]{pgfstrokecolor}{rgb}{0.1333,0.5451,0.1333}M_{t}^{sf}$ $=M_{t}^{{}^{\prime}}-\lambda\cdot\sigma(\max(\log M_{t}^{{}^{\prime}},\log{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{usf}}))$
@@
Datasets
We evaluate SafeInfer on five existing datasets – DangerousQA (Shaikh et al. 2023), AdvBench (Zou et al. 2023b), HEx-PHI (Qi et al. 2023a), NicheHazardQA (Hazra et al. 2024), and TechHazardQA (Banerjee et al. 2024). Further, we propose a new safety dataset based on the list of violated policies identified by Meta (Qi et al. 2023a). We describe each of these datasets in detail below.
@@
HarmEval: To further evaluate our safety strategy SafeInfer across a broad spectrum of potential hazards, we developed HarmEval, a dataset based on prohibited scenarios listed in OpenAI and Meta’s usage policies, as cited in (Qi et al. 2023b). HarmEval categorizes risks into 11 main categories (see Figure 3), resulting in approximately $\sim$550 crafted harmful queries. We employed a two-step verification process for these queries. First, we used GPT-4 to classify the queries as harmful or non-harmful [^2] Prompt: Determine whether the question adheres to ethical standards or if there are certain contents that raise ethical concerns. Please answer only in either ‘Yes’ or ‘No’.. Following this, we utilized the Perspective API [^3] https://perspectiveapi.com/ to calculate the toxicity score, selecting only those queries with a toxicity score of 0.7 or higher. The dataset will be released upon the acceptance of our work.
@@
GCG AutoDAN PAIR DeepInception GPTFuzzer
AdvBench
Base Model 0.37 0.44 0.52 0.29 0.29
SafeDecoding 0.13 0.09 0.10 0.08 0.05
SafeInfer 0.07 0.04 0.02 0.01 0
HarmEval
Base Model 0.48 0.53 0.68 0.46 0.51
SafeDecoding 0.22 0.17 0.12 0.09 0.14
SafeInfer 0.02 0 0.01 0 0.02
Table 4: ASR of harmful responses for popular jailbreak methods for Llama-2.
@@
ARC OKTest GSM8K
Llama-2 Mistral Llama-2 Mistral Llama-2 Mistral Llama-2 Mistral Llama-2 Mistral Llama-2 Mistral
Base model 17.83 5.22 46.90 62.00 0.298, 0.451 0.501, 0.656 0.416 0.525 0.14 0.08 22.29 51.9
SafeDecoding 80.30 - 45.70 - 0.376, 0.518 - 0.399 - 0.10 - 21.98 -
SafeInfer 20.09 5.22 46.47 61.60 0.390, 0.582 0.531, 0.691 0.416 0.532 0.10 0.06 22.07 51.5
@@
Base models: In accordance with (Jain et al. 2023), we utilize base model backbones such as Llama2-7b-chat-hf (Touvron et al. 2023) and Mistral-7B-Instruct-v0.2 (Jiang et al. 2023).
Edited models: Previous research (Banerjee et al. 2024; Hazra et al. 2024) has observed that edited models can introduce hidden harms after updating the knowledge of the model (model editing). Therefore, our method has been evaluated on edited models with the Llama2-7b-chat-hf backbone. We employ a locate-and-edit model-based algorithm known as ROME (Meng et al. 2022). Our primary goal is to examine the impact of model editing on model safety, which is why we opted for a single edit algorithm (ROME) and a single model (Llama-2). For the most part, we utilize the default parameter values provided in paper (Hazra et al. 2024).
Prompting technique
@@
For simple prompts, we employed the vanilla strategy by directly asking the questions present in the datasets and expecting the model to generate responses. Recent studies by (Banerjee et al. 2024) have demonstrated that models can be ‘jailbroken’ by prompting them in an instruction-centric manner. This is followed by instruction-centric CoT prompts, which infuse unethical content more effectively into the generated responses. Inspired by this, we conduct experiment using instruction-centric and instruction-centric CoT prompts.
To assess the defense performance when a naive attacker directly inputs harmful queries to the language model, we utilized the six datasets mentioned previously. Detailed setups of these prompting techniques can be found in the Appendix.
Baselines
We evaluate our proposed method against the following safety alignment baselines following a decoding-based approach: SafeDecoding (Xu et al. 2024) and Self-CD (Shi et al. 2024) methods. Further, we directly use SA and sGDS as a standalone baselines to establish the effectiveness of the amalgamation of the two techniques.
SafeDecoding: SafeDecoding (Xu et al. 2024) is a safety decoding strategy used while responding to user queries. This approach is built upon the crucial observation that tokens representing safety warnings are often ranked high in probability, even when harmful content tokens are also prevalent. By selectively boosting the probability of these safety tokens and diminishing the likelihood of harmful sequences, SafeDecoding effectively counters the risks posed by jailbreak attacks. We show the results of the Llama2-7b model. Due to the lack of knowledge about the fine-tuning dataset used, we could not reproduce the results for Mistral-7b.
Self-CD: We also compare SafeInfer against Self-Contrastive Decoding (Self-CD) (Shi et al. 2024), which mitigates the issues of harmfulness as well as helpfulness. Self-CD is designed as a training-free and model-independent intervention, which attempts to amplify the difference in output token distributions when responding to questions with a safety prompt and without a safety prompt. The final next token distribution is determined by removing the over-attention from the model via contrastive decoding.
@@
Jailbreak methods
We examine five state-of-the-art jailbreak attacks, each representing a different category. Among these, GCG (Zou et al. 2023c) employs a gradient-based approach, while AutoDAN (Liu et al. 2024b) utilizes genetic algorithms and PAIR (Chao et al. 2024) utilizes an edit-based attack. In addition, we also explore DeepInception (Li et al. 2024) and GPTFuzzer (Yu et al. 2024) as key examples of empirical jailbreak attacks. To evaluate the effectiveness of our defenses against straightforward harmful query inputs, we use two datasets: Advbench and HarmEval.
Evaluation metric
We follow the methodology outlined by (Liu et al. 2024c) and utilize attack success rate (ASR) to evaluate the effectiveness of SafeInfer. ASR is defined as follows.
@@
Simulating GPT-4 as primary evaluator alongside secondary human evaluation: Manually evaluating each generated response for potential harm is both tedious and mentally taxing. To address this, inspired by several prior works (Qi et al. 2023b; Zheng et al. 2023), we employ GPT-4 as an automated evaluator to simulate human judges. We instruct GPT-4 to analyze and assess the ethical implications of responses generated by our test models. The specific instruction given to GPT-4 is: “Given the following text, analyze and assess its ethical implications. Determine whether the content adheres to ethical standards or if there are elements that raise ethical concerns.” GPT-4 is then tasked with providing a ‘YES’ or ‘NO’ answer regarding the presence of ethical concerns. To ensure the accuracy of GPT-4’s evaluations, we randomly sample 40% of the model-generated responses and obtain human judgments for comparison. Four experts with experience in working on AI safety partook in the evaluation. The average pairwise inter-annotator agreement (Cohen’s $\kappa$) is 0.94. Our analysis reveals a high concordance rate, with GPT-4’s judgments matching human judgments in $>93\%$ cases. Note that to manage computational costs, we select a stratified sample of up to 100 questions from of the each categories present in the NicheHazardQA, TechHazardQA, and HarmEval datasets. When fewer than 100 questions were available in a category, we use all available questions. We average the results from over all the categories. For other datasets – DangerousQA, AdvBench, and HEx-PHI – we select $\sim$200 stratified questions. For every dataset the selected questions are fed to the model, and the responses are evaluated for safety using GPT-4 and humans.
Obtaining the harmful model
We construct a small set of safe demonstrations, $D_{sf}$, from our proposed HarmEval dataset, consisting of approximately $|\mathsf{P}|$ = 100 prompts. Each prompt, $\mathsf{p}$, includes 10 contextual samples (harmful question-safe answer (see samples in Appendix)) and a query. Further, we use the HarmEval dataset to create $\mathbb{D}_{usf}$, a collection of harmful question-answer pairs. Following (Qi et al. 2023a), we select around $\sim$100 queries and their harmful responses to finetune a model with the same base model as $M_{b}$ and obtain the harmful model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{usf}$.
Utility and over-safety test
To evaluate the utility of the model after applying the proposed method, we conduct thorough evaluation on MMLU (5 shots) (Hendrycks et al. 2021) and TruthfulQA (Lin, Hilton, and Evans 2022). For testing over-safety, we use the framework used by (Röttger et al. 2024) where the LLM backbone generates three main types of responses on the XSTest dataset: (1) full compliance (2) full refusal (3) partial refusal. We only count responses classified as full compliance as the refusal rate to measure over-safety.
Results